Bob leaves Pittsburgh traveling west. Jim leaves from the same point two hours later and travels the same route at 18 miles per hour faster than bob. If Jim wants to catch up four hours after he leaves, how fast will he have to go?
Found 2 solutions by ikleyn, ankor@dixie-net.com:
Answer by ikleyn(52879)

.
Bob leaves Pittsburgh traveling west. Jim leaves from the same point two hours later and travels the same route
at 18 miles per hour faster than bob. If Jim wants to catch up four hours after he leaves, how fast will he have to go?
~~~~~~~~~~~~~~~


A B S U R D I S T   formulation.

Read the post and try to understand what is written there,  if you can.

..........................

After reading the post by  @ankor@dixie-net.com,  I can confirm again  (repeat one more time)  my opinion that
the problem,  as it is worded,  printed,  posted and presented,  is   A B S U R D I S T.

So, if you are planning to distribute it around,  you better throw it to the GARBAGE BOX,  which is the only right place for it.

Otherwise,  people will be laughing on it.


/\/\/\/\/\/\/\/

Do not forget to submit your "deepest THANKS" to me for my reaction and advises.


--------------------------

                To the managers of this project !


        Attention (!)     Attention (!)     Attention (!)


Last couple of days,  and especially today,  I observe a flow of  FAKE  PROBLEMS  posted to this forum.

Some list of them is below  (it does not cover all such posts).

https://www.algebra.com/algebra/homework/Triangles/Triangles.faq.question.1149289.html
https://www.algebra.com/algebra/homework/Probability-and-statistics/Probability-and-statistics.faq.question.1149284.html
https://www.algebra.com/algebra/homework/word/travel/Travel_Word_Problems.faq.question.1149270.html
https://www.algebra.com/algebra/homework/Probability-and-statistics/Probability-and-statistics.faq.question.1149083.html


I ask you to fire immediately the person who produced these posts,  since he  (or she)  is not able to provide the right input,
and to replace him  (or her)  by a better prepared and adequate person.

Thank you.



Answer by ankor@dixie-net.com(22740)

Bob leaves Pittsburgh traveling west.
Jim leaves from the same point two hours later and travels the same route at 18 miles per hour faster than bob.
If Jim wants to catch up four hours after he leaves, how fast will he have to go?
;
let s = Jim's speed
then
(s-18) = Bob's speed
:
When J catch's B it will be 6 hrs after B departed (2 + 4)
They will have traveled the same distance
Write a distance equation, dist = time * speed
6(s-18) = 4s
6s - 108 = 4s
6s - 4s = 108
2s = 108
s = 108/2
s = 54 mph
:
Check find the actual dist each traveled B's speed 54 - 18 = 36 mph
6 * 36 = 216 mi
4 * 54 = 216